lim (e^x +e^-x+2cosx -4)x^4
x-0

Grade:12

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

L = \lim_{x\rightarrow 0}\frac{e^{x}+e^{-x}+2cosx-4}{x^{4}}
It is zero by zero from. Use L’Hospital rule
L = \lim_{x\rightarrow 0}\frac{e^{x}-e^{-x}-2sinx}{4x^{3}}
L = \lim_{x\rightarrow 0}\frac{e^{x}+e^{-x}-2cosx}{4.3x^{2}}
L = \lim_{x\rightarrow 0}\frac{e^{x}-e^{-x}+2sinx}{4.3.2x}
L = \lim_{x\rightarrow 0}\frac{e^{x}+e^{-x}+2cosx}{4.3.2.1}
L = \frac{1+1+2}{4.3.2.1}
L = \frac{1}{6}

Abhishek Pachauri
9 years ago
its very simple.write all the expressionss.and get 1/6